### v1

# A Framework for Quantum Advantage

Abstract
As quantum computing approaches the threshold where certain tasks demonstrably outpace their classical machines, the need for a precise, clear, consensus-driven definition of quantum advantage becomes essential. Rapid progress in the field has blurred this term across companies, architectures, and application domains. Here, we aim to articulate an operational definition for quantum advantage that is both platform-agnostic and empirically verifiable. Building on this framework, we highlight the algorithmic families most likely to achieve early advantage. Finally, we outline our vision for the near future, in which quantum computers enhance existing high-performance computing platforms, enabling new frontiers in chemistry, materials discovery, optimization, and beyond.
I Introduction
Quantum computing has long held promise in solving computational problems that are intractable for classical computers [1, 2, 3, 4, 5, 6, 7]. With the astonishing improvement in hardware, software and quantum algorithms, this promise feels closer than ever before, e.g., as seen in several demonstrations that have shown how quantum processors are now capable of performing accurate computations at a scale that defies brute-force classical simulation of quantum circuits [8, 9]. As these capabilities continue to mature, a central question has come to the forefront: When will quantum computers deliver a tangible advantage in solving practical computational tasks?
This milestone, commonly referred to as quantum advantage, represents a pivotal threshold for the field. It serves not only as a benchmark for quantum technological progress, but also as a guiding framework for allocating research investments and identifying application domains with near-term impact potential. Establishing criteria for when and how quantum advantage is achieved is therefore essential for advancing both our scientific understanding and the deployment of quantum technologies in real-world settings.
In this work, we outline the general definition of quantum advantage and the criteria that are necessary to claim it, as well as the necessary elements to achieve it: methods of running accurate quantum circuits, HPC + quantum infrastructure, and performant quantum hardware. It is also crucial to focus our attention on algorithms that we know can be run reliably and have verifiably correct outputs, thus, we also outline the collection of algorithms that fit this description. Once quantum advantage has been attained in experimental settings, we envision that quantum computers will start enhancing existing HPC platforms, unlocking new frontiers in chemistry, materials science, optimization, and beyond.
II Definition of Quantum Advantage
Quantum advantage denotes the execution of an information-processing task on quantum hardware that satisfies two essential criteria: i) the correctness of the output can be rigorously validated, and ii) it is performed with a demonstrably superior efficiency, cost-effectiveness, or accuracy than is attainable with classical computation alone. Although the concept appears straightforward, these criteria remain insufficiently articulated in the current literature- e.g. in order to have confidence in the correctness of the output produced by a quantum computer, one must demonstrably validate that the quantum computation is trustworthy. As we argue in the manuscript, this is primarily achieved through rigorous error bars as can be obtained from error correction or formally proven error mitigation methods. The result can also be validated if it is efficiently verifiable on a classical computer. Given the accelerated pace of progress in the field, a consensus definition of quantum advantage is needed to enable rigorous evaluations of forthcoming claims.
III Problems poised for quantum Advantage
In this section, we discuss the family of problems that are strong candidates to reach quantum advantage. Strong claims of advantage need to rely, at least, on one of two things: verifiability of the result, or trust in the underlying quantum hardware, algorithms and methods executing the computation. In the context of current quantum processors, this could present a substantial challenge since, while certain problems allow for direct classical verification of quantum outputs, many relevant problems of interest do not [10]. In these cases, confidence must be established through the indirect certification of the quantum computation itself. A fault tolerant quantum computer (FTQC) is one way to make this certification, however, rigorous methods are emerging where the full overhead of FTQC is not needed. This is a necessary precursor for quantum advantage, and has been deemed in the literature as quantum utility [8]: the ability of quantum devices to perform accurate computations at scales beyond exact brute-force classical simulations. Provided that quantum utility is established, the focus shifts to identifying computational tasks and algorithms where quantum solutions provide evidence supporting the claim of advantage.
In what follows, we discuss three key families of computational problems that currently stand out as promising avenues for realizing quantum advantage: i) problems that are solved via sampling algorithms, ii) problems that are verifiable via the variational principle, and iii) problems whose solution reduces to calculating expectation values of observables. We examine each of these approaches in greater detail, with the aim of elucidating their respective potential for demonstrating quantum advantage.
Sampling problems
Sampling algorithms use the information contained in a set of samples from probability distributions defined by quantum circuits. According to our current understanding of computational complexity, these algorithms are well poised to solve problems that are difficult for classical computers, such as integer factorization using Shor’s algorithm [2]. Here, however, the apparent need for error correction and fault tolerance represents a major challenge that we do not anticipate to be solved in the short term.
Another task that falls into the sampling category is random circuit sampling (RCS) [11]. RCS could offer a potential candidate for quantum advantage, but verifying the correctness of the solutions to these problems is classically intractable at scale. A natural way to certify such computations is to ensure that the method of producing the samples is free (or nearly free) of errors. To date, the only way we know how to achieve this is through error correction, which is why several experimental demonstrations [12, 13, 14] that claim quantum advantage have, in our view, not yet achieved this result. Attempts to explore sampling with quantum computers in the analog mode using neutral atoms [15] or superconducting qubits in the XY model [16], are promising from a physics perspective, but raise similar concerns regarding verification and robustness. In short, these demonstrations fail to certify the accuracy of the computation, which negates a fair comparison between quantum and classical methods [17]. Overall, given the lack of efficient verification, we do not believe that random circuit sampling is a fully satisfactory route toward quantum advantage.
However, recent theoretical work [18] has introduced a promising variant of random circuit sampling (RCS) based on peaked random circuits. Such circuits are engineered to produce a specific bitstring with non-negligible probability, while maintaining random-like behavior otherwise. This structure opens new avenues for demonstrating verifiable quantum advantage, as the “peakedness” of the distribution can be detected by analyzing sampled outputs. Although recent results show that peakedness may, in some scenarios, be simulated in quasi-polynomial time [19], this insight also enhances our understanding of when and how quantum circuits exhibit a classically tractable structure. Furthermore, the possibility that classical algorithms might directly identify peakedness, without full simulation, offers valuable feedback for designing more robust quantum advantage protocols. Overall, peaked random circuits provide fertile ground for exploring new complexity-theoretic frontiers towards verifiable quantum advantage, but will need more rigorous determination on hardness to be clear advantage candidates.
Variational principle type of problems
A family of computational problems that are promising candidates for achieving quantum advantage is those that obey the variational principle. One such example is finding the ground-state energies in certain physical systems. Computing ground-state energies is an important application at the root of many computational challenges in condensed matter and high-energy physics, chemistry, and materials science. The ground-state problem has an important implication in the search for quantum advantage. First, if the algorithm used satisfies the variational principle, approximate solutions can be classically ranked by accuracy, according to the energy approximations they produce. As a consequence, if a quantum solution has a better accuracy or lower energy, that would indicate a possible verifiable quantum advantage.
Verifiability can occur at different levels depending on the nature of the algorithm. One can have variational quantum solutions that come with an uncertainty error, without it, or - for some algorithms - the entire solution proposed by the quantum computer can be classically representable. In the following, we discuss a few examples of quantum algorithms that have a range of verifiability levels.
The quantum phase estimation algorithm (QPE) [20] is one of the first algorithms proposed for the ground-state problem. Provided with a good reference state with a polynomial overlap with the true ground state, QPE can solve the ground state problem efficiently in polynomial time. The solution from QPE algorithm comes with a precision error which is exponentially small in the size of its quantum auxiliary register, and its verifiability hinges on the trust that the quantum circuit is performed correctly and other approximations - such as those accompanying the Hamiltonian simulation part - are controlled. However, the quantum circuits that are necessary for the phase estimation algorithm are very deep, and they are generally thought to require fault-tolerant quantum computers.
In the last decade, variational quantum eigensolvers (VQEs) [21, 22] and their adaptive versions [23] have been proposed as a low-depth alternative for ground-state problems, which can be executed on pre-fault-tolerant quantum devices. They can be seen as quantum equivalent of variational quantum Monte Carlo methods [24]. Compared to phase estimation, they do not have convergence guarantees, and are not directly tied to the quality of the initial reference state. Although the approximate estimations of the ground-state energies produced by VQE are variational, they come with an uncertainty related to the statistical nature of the energy estimation. This uncertainty allows the approximate solution to underestimate the ground state energy, breaking a strict ranking of the solution accuracy. However, one can take into account these uncertainties and introduce quality metrics that incorporate them, such as the V-score defined in [25]. While still constituting a good option for lattice models, low-depth VQE is not a good choice for utility-scale chemistry systems, since the number of measurements necessary to estimate molecular energy to useful precisions is prohibitive [26, 27].
The Krylov quantum diagonalization (KQD) [28], a quantum equivalent of classical algorithms based on power methods, bridges QPE and VQE, since it inherits convergence properties of QPE, and the variational behavior and statistical uncertainties of VQE. It removes the need to design a variational ansatz, and hinges on quantum circuits that simulate real-time evolutions. The verifiability of its solutions can be treated on the same level of VQE, and the solution accuracy can be ranked accordingly. Still, KQD cannot deal with large chemistry problems as a result of the same measurement problem that affects VQE.
Recently, new algorithms designed to be executed on quantum-centric supercomputing platforms [9, 29] have overcome different issues that were affecting pre-fault-tolerant ground-state methods. Instead of estimating expectation values, as in VQE or KQD, the sample-based quantum diagonalization (SQD) uses measurements from a quantum ansatz state to identify a subspace onto which the original Hamiltonian can be projected. The projected Hamiltonian has a much lower dimension than the original and can be diagonalized using distributed classical computing. Remarkably, SQD has already been enabled to use quantum data to estimate energies for systems such as the iron sulfide cluster 2Fe-2S, matching classical state-of-the-art results, and tackled larger 4Fe-4S clusters beyond exact diagonalization [9]. The fact that the energies are estimated on a classical computer at machine precision allows one to estimate molecular energies that exceed what is achievable with VQE, at the same time dealing with errors from a quantum computer at classical machine precision, effectively removing the impact of useless quantum samples at the cost of a classical overhead.
The SQD approach can be further combined with KQD, by classically projecting a target Hamiltonian on a subspace of samples from real-time evolution circuits, in the sample-based Krylov quantum diagonalization (SKQD) workflow [30]. The benefit of incorporating SQD and KQD is twofold: first, it comes with convergence guarantees similar to phase estimation [31], and second, it does not require the design of a quantum circuit ansatz.
All the sample-based methods for ground states, both SQD and SKQD, achieve the highest degree of verifiability. The fact that the approximate ground states produced by these methods can be stored in classical memory allows for a high-precision ranking of their accuracy, without the need to factor in uncertainties coming from statistical noise. Quantum advantage obtained with these methods would therefore be classically certifiable and reproducible.
Quantum computers in the analog mode also allow one to generate bitstring distributions for regimes that can be challenging for classical methods. This includes regimes close to criticality or frustrated systems. Such examples are potential candidates for quantum-centric approaches such as SQD or SKQD.
Similarly to sample-based methods for quantum systems, the quantum approximate optimization algorithm (QAOA) outputs candidate solutions for combinatorial problems [32] that are classically verifiable. While we do not expect a general quantum speed-up for most optimization problems, we see potential for finding better solutions than classical solvers for specific problem instances with QAOA and other quantum heuristic approaches. To track progress towards potential quantum advantage and drive a fair and systematic benchmarking effort, a quantum optimization benchmark library has been introduced, with a curated set of 10 combinatorial optimization problem classes, corresponding problem instances, and well-defined reporting metrics [33]. Along similar lines, Cazals et al. [34] have identified hard instances that can be natively suited for neutral-atom-based analog quantum computers. Furthermore, neutral atoms arranged in amorphous arrays have been proven to have a spin-glass phase [35], which is expected to be hard to sample from by classical means.
Crucially, for many problem classes, instances already become challenging for state-of-the-art classical solvers at moderate problem sizes for known sparse model formulations. This indicates a strong compatibility with near-term quantum hardware. Nevertheless, based on the comparative strength of classical optimization methods, we believe that the most promising path forward is to target individual problem classes and even specific problem instances.
Expectation values of observables
The calculation of expectation values of observables of short-depth quantum circuits is an essential task for several near-term quantum algorithms targeting a range of computational problems: from estimating ground-state energies of molecules [21], measuring kernels in machine learning [36], to simulating the time evolution of spin correlation functions [8]. In the absence of fault tolerance, these estimations have also been shown to be biased by noise in early implementations of variational algorithms [22] with modest circuit depths. However, the invention of error mitigation methods, detailed later in Section IV, has provided a path to suppressing [37, 38] and even canceling [37] the effect of noise on obeservable estimation, albeit at exponential sampling cost. However, improvements in hardware error rates and sampling rates have scaled these methods to qubit and gate counts that are beyond brute-force classical computation [8] and could even outperform state-of-the-art classical approximations.
Simulating time evolution is a classic example of measuring expectation values of short-depth quantum circuits and a strong candidate for quantum advantage, that is important to consider from the lens of verifiability. Models with conserved quantities offer some verification of the quantum computation, but in the most general case, verifiability can be a challenge at scales beyond exact diagonalization. At such scales, as we will elaborate more in Section IV, it is particularly crucial to have error mitigation methods that can provide expectation values with rigorous error bars, to place trust in the result of the quantum computation.
Simulating time evolution is also particularly well suited for analog quantum simulators, and has already been studied with these systems at scales inaccessible to exact diagonalization, particularly for specific classes of Hamiltonians [39]. While controlling and characterizing errors in these systems remains an open challenge, recent theoretical results suggest that certain physically meaningful observables (e.g., local) can exhibit robustness to both deterministic and stochastic perturbations, with error bounds that may not scale with system size [40, 41, 42]. These insights, together with ongoing developments in analog error mitigation [43, 44], point to promising paths to enhance the reliability of analog quantum simulation.
Based on these algorithmic groups, several application domains are widely considered the most promising areas that could quickly take advantage of quantum advantage. Quantum simulations of chemistry and materials science [9, 45, 46, 47, 48] are widely regarded as one of the most promising near-term applications. These problems are often intractable for classical computers due to the exponential complexity of simulating quantum systems with high accuracy. Quantum computers, in contrast, can efficiently represent other quantum systems, and quantum algorithms exist to simulate both static and dynamical properties of quantum systems that have no efficient classical counterpart and can, in principle, lead to quantum advantages.
IV Elements for Advantage
In what follows, we discuss the necessary elements critical to reach and certify quantum advantage in the problems outlined above. These elements include methods capable of running accurate circuits. Such accuracy can be established through error bars, obtained from error correction, or rigorous error mitigation and detection methods. Finally, we discuss the role that quantum-centric supercomputing infrastructure, and performant quantum hardware will play towards achieving quantum advantage.
IV.1 Methods for running accurate circuits
1.
Error correction Quantum error correction [49] (QEC) refers to a set of techniques designed to protect quantum information from decoherence, gate imperfections, and measurement errors—limitations intrinsic to any practical quantum computing platform. In QEC, information is encoded into logical qubits, fault-tolerant constructs built from multiple physical qubits, such that errors on a subset of physical qubits can be detected and corrected without collapsing the encoded state. QEC lays the groundwork for fault-tolerant quantum computation (FTQC), where quantum algorithms can be performed in a manner that is resilient to noise, provided that the physical error rates remain below a threshold. Among various QEC codes, the surface code [50] has been a leading candidate, for it is particularly suited for 2D architectures like superconducting qubits, and it offers high error thresholds close to 1% per operation [51]. However, this code requires a large overhead—on the order of thousands of physical qubits per logical qubit for realistic noise models—posing a scalability challenge. To address this, quantum low-density parity-check (qLDPC) codes have emerged as a compelling alternative. These codes can offer similar levels of error suppression with a 10x reduction in qubit overhead [52]. This makes them especially attractive for implementing deep-circuit algorithms such as Shor’s algorithm [2] and quantum phase estimation (QPE) [53], which are potential candidates for achieving quantum advantage in areas such as quantum chemistry and integer factorization.
This correctable nature of fault-tolerant circuits ensures the accuracy of quantum computations. However, given the substantial qubit and operational overhead required for encoding, syndrome extraction, and correction, the implementation of full QEC is not expected to be viable in the near term, although recent algorithmic improvements could help change this paradigm [54]. As such, early demonstrations of quantum advantage will likely rely on other schemes such as error mitigation and error detection strategies.
2.
Error mitigation Quantum error mitigation (QEM) refers to the set of techniques that aim to reduce, and sometimes eliminate, the error bias in expectation values caused by noise in quantum circuits [37, 55]. Unlike QEC, QEM techniques use classical post-processing of the noisy output from an ensemble of circuit runs to “mitigate” the effect of noise. Common QEM approaches include readout error mitigation to correct for measurement assignment errors [56, 57], and zero-noise extrapolation (ZNE), in which the noise strength is amplified, with the ultimate goal of “extrapolating” results toward the zero-noise limit [38, 58, 8]. Even in the absence of fault tolerance, some QEM methods can provide accurate estimates of expectation values measured off short-depth quantum circuits, with rigorous error bounds [59, 60]. However, such methods incur significant sampling overhead, often exponential in noise strength, due to the amplification of variance introduced by the quasi-probabilistic weighting. This is the case for probabilistic error cancellation (PEC), an error mitigation method that effectively inverts well-characterized noise channels [60]. The runtime cost for PEC roughly scales as for a -qubit circuit of depth , and hardware two-qubit depolarizing error rates [60]. Despite the exponential scaling, as hardware error rates drop, these methods can provide speed-ups even over classical methods with polynomial runtime scaling [61], at the 100-qubit scale with thousand’s of two-qubit gates. We project that further improvements in hardware error rates will drive the reach of error mitigation to increasingly challenging circuits/observables and circuit depths towards the realization of quantum advantage [62, 63, 64].
The reach of these methods can be further extended with classical HPC, than can further reduce the overhead for error mitigation error mitigation in exchange for classical runtime and potential bias, while leveraging the latest advances in classical simulation. This implies that at a given error rate, one can mitigate upon even larger circuit volumes. We have already seen this using tensor network error mitigation (TEM) [65], shaded light cones for PEC [66], and multi-product formulas (MPF) for algorithmic error mitigation of trotterized time evolution [67]. With these advances, we are confident that we will soon reach a point where quantum methods will outperform existing classical-only methods.
3.
Error detection Lastly, we see promise in post-selected quantum error detection methods. A simple way to view the difference between error correction and error mitigation is to see them as a trade-off between qubits (quantum resources) and sampling (classical resources). Error detection offers a more intermediate goalpost: use extra qubits to detect but not correct errors. This includes error-correcting codes with post-selection, which are also subsumed under error detection. Error detection can have a better (though still exponential) sampling overhead than other error mitigation methods that invert the noise, at the expense of using some extra qubits (though far fewer than full error correction). Post-selected error detection is the process of detecting when a run of the circuit (a ”shot”) has been corrupted by an error, and post-selecting for good shots. The post-selection rate falls exponentially with the circuit volume and the underlying physical noise rate. Unlike other error mitigation approaches, this benefits from an extremely fast sampling speed. In addition to its better sampling overhead, error detection has another appeal, namely that it works on a single-shot basis, not solely on expectation values of observables. Effectively, we slow down the quantum computer but make it produce higher quality samples. This makes error detection compatible with any other error mitigation approach. The combination of better sampling overhead and its single-shot nature makes error detection a promising candidate for achieving finite (but not asymptotic) quantum advantage.
Various methods have been proposed for error detection. The most obvious approach is to use an error detection code such as the [[4, 2, 2]] code or the [[8, 3, 2]] code [68, 69, 70, 71, 72, 73]. However, other alternatives exist for detecting errors not based on encoding of qubits but based on verifying symmetries in whole circuits [74]. For example, we might expect that a particular circuit preserves particle number, and any measurement that reveals a violation of this expectation could indicate an error [75]. In recent years another approach based on Pauli checks has been introduced which can detect errors in Clifford-dominated circuits by non-destructively measuring Pauli operators around the circuit and verifying that the Paulis were correctly mapped to each other through the Clifford [76]. Each measurement acts as a check and detects some errors.
Since the process of error detection itself has extra overhead in the terms of qubits and gates, it can introduce new errors in the circuit. It is therefore crucial to design low-overhead error detecting circuits so that their net effect is the removal and not increase of errors. This was recently demonstrated in the framework of spacetime codes, with error detection in large graph states comprising up to 50 qubits and 2500 gates, and also creating the largest GHZ states on record of 100+ qubits [77].
IV.2 Quantum-centric supercomputing infrastructure
A widespread misconception frames quantum advantage as a contest between quantum and classical machines. In practice, however, the relevant benchmark is not quantum systems in isolation, but hybrid quantum–classical systems that combine quantum processing units (QPUs) with classical supercomputers. We expect QPUs to enhance CPUs and GPUs capabilities within a heterogeneous computing model, tackling certain computational problems more efficiently. This emerging paradigm, which we call quantum-centric supercomputing (QCSC in short), positions classical processors as indispensable throughout the quantum computational workflow—handling system control, calibration, input processing such as circuit optimization, and postprocessing of quantum hardware outputs [37, 78]. The objective is therefore not to replace classical computing, but to demonstrate that an integrated hybrid workflow, with quantum computing at its core, can surpass what classical computation alone can achieve.
Realizing near-term quantum advantage, while paving a direct path to large-scale, fault-tolerant machines projected for the latter half of this decade, demands a systematic extension of the HPC architecture so that modular QPUs can be orchestrated as additional resources. At the hardware level, we envision QPUs as integral components of quantum-centric supercomputers—hybrid systems where QPUs act as specialized co-processors integrated with classical compute elements like CPUs and GPUs at different levels of tightness to enable seamless deployment of hybrid quantum-classical algorithms. Pasqal and IBM quantum computers are already being deployed in high-performance computing centers alongside classical resources[9, 79]. And both companies recently announced the release of a SLURM plugin developed with the HPC community aimed at making the integration of QPUs in HPC centers standard and seamless [80].
This set of tools and capabilities will be instrumental in heterogeneous algorithm development through a modular approach in which the different algorithmic pieces can be isolated and benchmarked for performance. The output of this process should be aimed towards reconciling the innate parallelism of modern HPC programming models with the way quantum workloads are defined, optimized, and deployed. Algorithm optimization should leverage workflow management capabilities (e.g. leveraging real-time sampling in variational approaches like SQD for more efficient or faster time-to-solution outcomes).
Extending the tools and infrastructure already familiar to the HPC community is essential to accelerating the adoption of quantum computing across the broader scientific and industrial ecosystem.
IV.3 Performant quantum hardware
Achieving quantum advantage in the short term is fundamentally constrained by the performance of the underlying hardware. In particular, the number of qubits, the achievable algorithmic depth, the fidelity of qubit operations, and the efficiency of quantum–classical orchestration, collectively set the ceiling for any credible claim of advantage. Among the various quantum architectures under development, superconducting quantum processors [81] and neutral atoms [82] stand out as particularly well positioned to reach this milestone, as both have recently surpassed the 100-qubit threshold and demonstrate promising progress across all key performance dimensions. Other platforms, such as trapped-ion quantum processors [83], offer a compelling alternative path to quantum computation. These systems have been deployed in cloud-accessible environments and are already being integrated into early-stage hybrid quantum-classical workflows. While our focus in this work will remain on superconducting circuits and neutral atom systems, the broader landscape of quantum hardware continues to evolve rapidly, with multiple modalities contributing to progress toward quantum advantage.
Superconducting qubits
Superconducting quantum processors are among the most technologically advanced platforms to date. They offer high-fidelity, high-speed gate operations, modularity, and compatibility with integrated control electronics.
IBM’s quantum processors, for example, employ transmon qubits arranged in modular, scalable architectures [84, 85]. The current 156-qubit Heron QPU achieves 250+ s relaxation times, error per layer gate (EPLG) of [85], and supports circuit layer operations per second (CLOPS) above 250,000. [9, 86]. Real-time classical controllers enable dynamic circuits and mid-circuit measurement and feed-forward, allowing for error detection and mitigation protocols.
Execution of hybrid quantum-classical workloads is supported by Qiskit, which provides a cloud-based runtime environment for low-latency quantum-circuit execution with classical feedback [87]. This framework integrates seamlessly in a serverless environment, orchestrates distributed hybrid quantum–classical tasks across containerized back-ends, and provides tools for domain-specific compilation, optimization, and error mitigation routines. Together, these components reduce overhead in iterative algorithms and enables a high-performance development environment that extends the practical reach of near-term devices toward practical demonstrations of quantum advantage.
Neutral atoms
Neutral-atom quantum computers are rapidly emerging as strong candidates for demonstrating quantum advantage, owing to their scalability, flexible connectivity, and ability to operate in both analog and digital modes [88]. These platforms trap individual atoms in programmable optical tweezers and control their interactions via Rydberg excitation, enabling the realization of both Ising-type analog dynamics and gate-based digital quantum circuits.
As a proof of this, Pasqal’s neutral-atom platform has already demonstrated 100+ qubits operating in analog mode and is expected to reach 250 qubits by 2026. In the analog regime, neutral atoms are naturally suited to simulate spin models, quantum dynamics and solve optimization problems, with demonstrated capabilities for addressing complex many-body systems. Pasqal QPUs are already being deployed alongside classical HPC resources at Tier-1 centers like GENCI and Jülich, where they can be accessed through standard schedulers and integrated into hybrid CPU/GPU/QPU workflows.
Neutral atoms can also support digital, gate-based quantum computing, aiming at early demonstrations of fault tolerance. This provides a clear path from near-term application development to more robust, general-purpose capabilities.
V Conclusion and future outlook
We maintain that credible evidence of quantum advantage is likely to emerge in one of the areas highlighted here within the next two years, contingent on sustained coordination between the high-performance-computing and quantum-computing communities, and most likely in the areas we have highlighted here.
Once these claims begin to emerge, they must be assessed on a common basis according to the two criteria articulated in this article. Furthermore, quantum advantage is unlikely to materialize as a single, definitive milestone. Instead, it should be understood as a sequence of increasingly robust demonstrations. Any claim of quantum advantage should be framed as a falsifiable scientific hypothesis, one that invites scrutiny, reproduction, and challenge from the broader community. A cautious iterative and interpretive stance is therefore essential until state-of-the-art classical methods have had adequate opportunity to challenge the claim. However, should a classical approach later supersede a previously reported quantum advantage benchmark, the outcome does not signify failure; rather, it exemplifies constructive behavior that drives progress in both classical and quantum computation.
Finally, we envision that the following best practices will be necessary to underpin credible advantage claims: (i) definition of standardized benchmark problems jointly with classical experts to ensure relevance and fairness, (ii) publication of detailed methodologies and datasets, sufficient to enable reproducibility of the benchmarks, and (iii) maintenance of an open-access leaderboards to track evolving performance. These practices will help ensure that quantum advantage claims are not only credible, but also constructive for the broader computational sciences.
Acknowledgements.
The authors thank the IBM Quantum and Pasqal teams for fruitful discussions and comments on the manuscript.References
- Feynman [1982] R. P. Feynman, International Journal of Theoretical Physics 21, 467 (1982).
- Shor [1997] P. W. Shor, SIAM Journal on Computing 26, 1484–1509 (1997).
- Deutsch [1985] D. Deutsch, Proceedings of the Royal Society A 400, 97–117 (1985).
- Harrow et al. [2009] A. W. Harrow, A. Hassidim, and S. Lloyd, Physical Review Letters 103 (2009).
- Lloyd [1996] S. Lloyd, Science 273, 1073–1078 (1996).
- Grover [1997] L. Grover, Physical Review Letters 79, 325–328 (1997).
- Preskill [2018] J. Preskill, Quantum 2, 79 (2018).
- Kim et al. [2023] Y. Kim, A. Eddins, S. Anand, K. X. Weia, E. van den Bergand Sami Rosenblatt, et al., Nature 618, 500 (2023).
- Robledo-Moreno et al. [2025] J. Robledo-Moreno, M. Motta, H. Haas, A. Javadi-Abhari, P. Jurcevic, et al., Science Advances 11, eadu9991 (2025).
- Eisert et al. [2020] J. Eisert, D. Hangleiter, N. Walk, I. Roth, D. Markham, et al., Nature Reviews Physics 2, 382 (2020).
- Aaronson and Chen [2016] S. Aaronson and L. Chen, (2016), arXiv:1612.05903 .
- Arute et al. [2019] F. Arute, K. Arya, R. Babbush, D. Bacon, J. C. Bardin, et al., Nature 574, 505 (2019).
- Villalonga et al. [2024] B. Villalonga, X. Mi, S. Mandra, A. Bengtsson, P. Klimov, Z. Chen, S. Hong, C. Erickson, I. Drozdov, et al., Nature 634, 328 (2024).
- DeCross et al. [2025] M. DeCross, R. Haghshenas, M. Liu, E. Rinaldi, J. Gray, et al., Phys. Rev. X 15, 021052 (2025).
- Shaw et al. [2024] A. L. Shaw, Z. Chen, J. Choi, D. K. Mark, P. Scholl, R. Finkelstein, A. Elben, S. Choi, and M. Endres, Nature 628, 71 (2024).
- Andersen et al. [2025] T. I. Andersen, N. Astrakhantsev, A. H. Karamlou, J. Berndtsson, J. Motruk, et al., Nature 638, 79 (2025).
- Kechedzhi et al. [2024] K. Kechedzhi, S. V. Isakov, S. Mandrà, B. Villalonga, X. Mi, S. Boixo, and V. Smelyanskiy, Future Generation Computer Systems 153, 431 (2024).
- Aaronson and Zhang [2024] S. Aaronson and Y. Zhang, (2024), arXiv:2404.14493 .
- Bravyi et al. [2023] S. Bravyi, D. Gosset, and Y. Liu, (2023), arXiv:2309.08405 .
- Kitaev [1995a] A. Y. Kitaev, arXiv preprint quant-ph/9511026 (1995a).
- Peruzzo et al. [2014] A. Peruzzo, J. McClean, P. Shadbolt, M.-H. Yung, X.-Q. Zhou, et al., Nature communications 5, 4213 (2014).
- Kandala et al. [2017] A. Kandala, A. Mezzacapo, K. Temme, M. Takita, M. Brink, J. M. Chow, and J. M. Gambetta, Nature 549, 242 (2017).
- Grimsley et al. [2019] H. R. Grimsley, S. E. Economou, E. Barnes, and N. J. Mayhall, Nature communications 10, 3007 (2019).
- Ceperley et al. [1977] D. Ceperley, G. V. Chester, and M. H. Kalos, Physical Review B 16, 3081 (1977).
- Wu et al. [2024] D. Wu, R. Rossi, F. Vicentini, N. Astrakhantsev, F. Becca, et al., Science 386, 296 (2024).
- Wecker et al. [2015] D. Wecker, M. B. Hastings, and M. Troyer, Phys. Rev. A 92, 042303 (2015).
- Zhang et al. [2025] Y. Zhang, X. Zhang, J. Sun, H. Lin, Y. Huang, D. Lv, and X. Yuan, Wiley Interdisciplinary Reviews: Computational Molecular Science 15, e70020 (2025).
- Yoshioka et al. [2024] N. Yoshioka, M. Amico, W. Kirby, P. Jurcevic, A. Dutt, et al., (2024), arXiv:2407.14431 .
- Kanno et al. [2023] K. Kanno, M. Kohda, R. Imai, S. Koh, K. Mitarai, W. Mizukami, and Y. O. Nakagawa, (2023), arXiv:2302.11320 [quant-ph] .
- Yu et al. [2025] J. Yu, J. R. Moreno, J. T. Iosue, L. Bertels, D. Claudino, et al., (2025), arXiv:2501.09702 .
- Kirby [2024] W. Kirby, Quantum 8, 1457 (2024).
- Farhi et al. [2014] E. Farhi, J. Goldstone, and S. Gutmann, (2014), arXiv:1411.4028 .
- Koch et al. [2025] T. Koch, D. E. B. Neira, Y. Chen, G. Cortiana, D. J. Egger, et al., (2025), arXiv:2504.03832 .
- Cazals et al. [2025] P. Cazals, A. François, L. Henriet, L. Leclerc, M. Marin, et al., (2025), arXiv:2502.04291 .
- Brodoloni et al. [2025] L. Brodoloni, J. Vovrosh, S. Julià-Farré, A. Dauphin, and S. Pilati, arXiv preprint arXiv:2505.05117 (2025).
- Havlíček et al. [2019] V. Havlíček, A. D. Córcoles, K. Temme, A. W. Harrow, A. Kandala, J. M. Chow, and J. M. Gambetta, Nature 567, 209 (2019).
- Temme et al. [2017a] K. Temme, S. Bravyi, and J. M. Gambetta, Phys. Rev. Lett. 119, 180509 (2017a).
- Kandala et al. [2019] A. Kandala, K. Temme, A. D. Córcoles, A. Mezzacapo, J. M. Chow, and J. M. Gambetta, Nature 567, 491 (2019).
- Choi et al. [2016] J.-y. Choi, S. Hild, J. Zeiher, P. Schauß, A. Rubio-Abadal, T. Yefsah, V. Khemani, D. A. Huse, I. Bloch, and C. Gross, Science 352, 1547 (2016), arXiv:1604.04178 [cond-mat.quant-gas] .
- Trivedi et al. [2024] R. Trivedi, A. Franco Rubio, and J. I. Cirac, Nature Communications 15, 6507 (2024).
- Cai et al. [2024] Y. Cai, Y. Tong, and J. Preskill, in 19th Conference on the Theory of Quantum Computation, Communication and Cryptography (TQC 2024), Vol. 310 (2024) pp. 2:1–2:15.
- Daley et al. [2022] A. J. Daley, I. Bloch, C. Kokail, S. Flannigan, N. Pearson, M. Troyer, and P. Zoller, Nature 607, 667 (2022).
- Guo et al. [2025] R.-C. Guo, Y. Gu, and D. E. Liu, npj Quantum Information 11, 14 (2025).
- Steckmann et al. [2025] T. Steckmann, D. Luo, Y.-X. Wang, S. R. Muleady, A. Seif, C. Monroe, M. J. Gullans, A. V. Gorshkov, O. Katz, and A. Schuckert, Error mitigation of shot-to-shot fluctuations in analog quantum simulators (2025), arXiv:2506.16509 .
- Georgescu et al. [2014] I. M. Georgescu, S. Ashhab, and F. Nori, Rev. Mod. Phys. 86, 153 (2014).
- Johnson et al. [2014] T. H. Johnson, S. R. Clark, and D. Jaksch, EPJ Quantum Technology 1, 10 (2014).
- Di Meglio et al. [2024] A. Di Meglio, K. Jansen, I. Tavernelli, E. Amico, M. Berta, et al., PRX Quantum 5, 037001 (2024).
- Alexeev et al. [2024] Y. Alexeev, M. Amsler, M. A. Barroca, S. Bassini, T. Battelle, et al., Future Generation Computer Systems 160, 666 (2024).
- Shor [1995] P. W. Shor, Phys. Rev. A 52, R2493 (1995).
- Fowler et al. [2012] A. G. Fowler, M. Mariantoni, J. M. Martinis, and A. N. Cleland, Phys. Rev. A 86, 032324 (2012).
- Benito et al. [2025] C. Benito, E. López, B. Peropadre, and A. Bermudez, Quantum 9, 1623 (2025).
- Bravyi et al. [2024] S. Bravyi, A. W. Cross, J. M. Gambetta, D. Maslov, P. Rall, and T. J. Yoder, Nature 627, 778 (2024).
- Kitaev [1995b] A. Y. Kitaev, (1995b), arXiv:quant-ph/9511026 [quant-ph] .
- Clinton et al. [2024] L. Clinton, T. S. Cubitt, R. Garcia-Patron, A. Montanaro, S. Stanisic, and M. Stroeks, (2024), arXiv:2410.21517 [quant-ph] .
- Li and Benjamin [2017] Y. Li and S. C. Benjamin, Phys. Rev. X 7, 021050 (2017).
- Nation et al. [2021] P. D. Nation, H. Kang, N. Sundaresan, and J. M. Gambetta, PRX Quantum 2, 040326 (2021).
- van den Berg et al. [2022] E. van den Berg, Z. K. Minev, and K. Temme, Phys. Rev. A 105, 032620 (2022).
- Giurgica-Tiron et al. [2020] T. Giurgica-Tiron, Y. Hindy, R. LaRose, A. Mari, and W. J. Zeng, in 2020 IEEE International Conference on Quantum Computing and Engineering (QCE) (IEEE, 2020) pp. 306–316.
- Temme et al. [2017b] K. Temme, S. Bravyi, and J. M. Gambetta, Phys. Rev. Lett. 119, 180509 (2017b).
- van den Berg et al. [2023a] E. van den Berg, Z. K. Minev, A. Kandala, and K. Temme, Nature Physics 19, 1116 (2023a).
- Schuster et al. [2024] T. Schuster, C. Yin, X. Gao, and N. Y. Yao, (2024), arXiv:2407.12768 .
- Temme et al. [2022] K. Temme, E. van den Berg, A. Kandala, and J. Gambetta, https://www.ibm.com/quantum/blog/gammabar-for-quantum-advantage (2022), blog post.
- Aharonov et al. [2025] D. Aharonov, O. Alberton, I. Arad, Y. Atia, E. Bairey, Z. Brakerski, I. Cohen, O. Golan, I. Gurwich, O. Kenneth, et al., arXiv preprint arXiv:2503.17243 (2025).
- Zimborás et al. [2025] Z. Zimborás, B. Koczor, Z. Holmes, E.-M. Borrelli, A. Gilyén, H.-Y. Huang, Z. Cai, A. Acín, L. Aolita, L. Banchi, et al., arXiv preprint arXiv:2501.05694 (2025).
- Fischer et al. [2024] L. E. Fischer, M. Leahy, A. Eddins, N. Keenan, D. Ferracin, et al., (2024), arXiv:2411.00765 .
- Eddins et al. [2024] A. Eddins, M. C. Tran, and P. Rall, (2024), arXiv:2409.04401 .
- Carrera Vazquez et al. [2023] A. Carrera Vazquez, D. J. Egger, D. Ochsner, and S. Woerner, Quantum 7, 1067 (2023).
- Harper and Flammia [2019] R. Harper and S. T. Flammia, Phys. Rev. Lett. 122, 080504 (2019).
- Gupta et al. [2024] R. S. Gupta, N. Sundaresan, T. Alexander, C. J. Wood, S. T. Merkel, et al., Nature 625, 259 (2024).
- Bluvstein et al. [2024] D. Bluvstein, S. J. Evered, A. A. Geim, S. H. Li, H. Zhou, et al., Nature 626, 58 (2024).
- Reichardt et al. [2024] B. W. Reichardt, D. Aasen, R. Chao, A. Chernoguzov, W. van Dam, et al., Demonstration of quantum computation and error correction with a tesseract code (2024), arXiv:2409.04628 .
- Reichardt et al. [2025] B. W. Reichardt, A. Paetznick, D. Aasen, I. Basov, J. M. Bello-Rivas, et al., Fault-tolerant quantum computation with a neutral atom processor (2025), arXiv:2411.11822 .
- Self et al. [2024] C. N. Self, M. Benedetti, and D. Amaro, Nature Physics 20, 219 (2024).
- Tsubouchi et al. [2025] K. Tsubouchi, Y. Mitsuhashi, R. Takagi, and N. Yoshioka, (2025), arXiv:2503.13114 .
- Bonet-Monroig et al. [2018] X. Bonet-Monroig, R. Sagastizabal, M. Singh, and T. E. O’Brien, Phys. Rev. A 98, 062339 (2018).
- van den Berg et al. [2023b] E. van den Berg, S. Bravyi, J. M. Gambetta, P. Jurcevic, D. Maslov, and K. Temme, Phys. Rev. Res. 5, 033193 (2023b).
- Martiel and Javadi-Abhari [2025] S. Martiel and A. Javadi-Abhari, (2025), arXiv:2504.15725 .
- McArdle et al. [2020] S. McArdle, S. Endo, A. Aspuru-Guzik, S. C. Benjamin, and X. Yuan, Reviews of Modern Physics 92 (2020).
- Liepuoniute et al. [2025] I. Liepuoniute, K. D. Doney, J. R. Moreno, J. A. Job, W. S. Friend, and G. O. Jones, Journal of Chemical Theory and Computation 21 (2025).
- Sitdikov et al. [2025] I. Sitdikov, M. E. Sahin, U. Bacher, A. Wennersteen, A. Damin, et al., (2025), arXiv:2506.10052 .
- Kjaergaard et al. [2020] M. Kjaergaard, M. E. Schwartz, J. Braumüller, P. Krantz, J. I.-J. Wang, S. Gustavsson, and W. D. Oliver, Annual Review of Condensed Matter Physics 11, 369 (2020).
- Saffman et al. [2010] M. Saffman, T. G. Walker, and K. Mølmer, Rev. Mod. Phys. 82, 2313 (2010).
- Bruzewicz et al. [2019] C. D. Bruzewicz, J. Chiaverini, R. McConnell, and J. M. Sage, Applied physics reviews 6 (2019).
- Bravyi et al. [2022] S. Bravyi, O. Dial, J. M. Gambetta, D. Gil, and Z. Nazario, Journal of Applied Physics 132 (2022).
- McKay et al. [2023] D. C. McKay, I. Hincks, E. J. Pritchett, M. Carroll, L. C. G. Govia, and S. T. Merkel, arXiV (2023).
- Community [2025] Q. Community, qiskit-device-benchmarking (2025), version 564164b, retrieved June 23, 2025.
- Javadi-Abhari et al. [2024] A. Javadi-Abhari, M. Treinish, K. Krsulich, C. J. Wood, J. Lishman, et al., (2024), arXiv:2405.08810 .
- Henriet et al. [2020] L. Henriet, L. Beguin, A. Signoles, T. Lahaye, A. Browaeys, G.-O. Reymond, and C. Jurczak, Quantum 4, 327 (2020).